OpenStudy (anonymous):

explain using linear approximations, why sin(x)≈x for x near 0

OpenStudy (anonymous):

perhaps you have seen that \[\lim_{x->0}\frac{\sin(x)}{x}=1\]that says the same thing

OpenStudy (anonymous):

or you can use "linear approximation" aka find the equation for the line tangent to \[\sin(x)\] at (0,0) the slope is the derivative evaluated at 0. the derivative of sine is cosine, and cosine of 0 = 1 so the line has slope 1

OpenStudy (anonymous):

and since it goes through (0,0) the equation for the line i s \[y=x\] so that is the linear approximation for \[\sin(x)\] near x = 0